OpenStudy (anonymous):

How would you find the equation of a line in standard form and slope-intercept form, that passes through (-1,5) and (1,9)?

OpenStudy (the_fizicx99):

M = y2-y1/x2-x1 plug in values

OpenStudy (the_fizicx99):

That'll give you your slope :p

OpenStudy (anonymous):

So the slope is 2?

OpenStudy (the_fizicx99):

Well, 9-5 = 4 1-(-1) = 2 so 4/2 = 2 yes

OpenStudy (the_fizicx99):

So naw that you have your slope, put it into y-y1=M(x-x1)

OpenStudy (anonymous):

The equation would be y=2x+7 then right?

OpenStudy (the_fizicx99):

So its: y-1=2(x-(-1) y-1=2x +2 +1 +1 y=2x + 3

OpenStudy (the_fizicx99):

How did you get 7 lol xD

OpenStudy (the_fizicx99):

@AlgebraOhNo

OpenStudy (anonymous):

Oh I put in a wrong value, for the y1 I put in 5 by accident

OpenStudy (the_fizicx99):

ah kay

OpenStudy (the_fizicx99):

Do you know what standard form is?

OpenStudy (anonymous):

Isn't that ax^2+bx+c?

OpenStudy (the_fizicx99):

Thats quadratic xD Ax + By = C

OpenStudy (the_fizicx99):

So how would you convert y = 2x +3 into Ax + By = C

OpenStudy (the_fizicx99):

@AlgebraOhNo

OpenStudy (the_fizicx99):

On a side note, do you know how to do quadratic? :O

OpenStudy (anonymous):

No not really

OpenStudy (the_fizicx99):

Not really as in for Ax + by = c or the quad thingy i asked lol <- descriptive

OpenStudy (anonymous):

The quadratic one

OpenStudy (the_fizicx99):

Oh lol kk, well that was irrelevant to this so, to convert y = 2x + 3 you would subtract 2x from both sides and get -2x + y = 3 .... in most cases you would multiply by -1 to get the x coefficient positive

OpenStudy (anonymous):

And that would be the standard form of the equation?

OpenStudy (the_fizicx99):

Idk some teachers allow the x coefficient to be negative, some require you to multiply is by -1 to make it positive

OpenStudy (anonymous):

So if I were to turn it positive it would be 2x-y=-3

OpenStudy (the_fizicx99):

If you'd like you just have to multiply every term by -1 -2x(-1) +y(-1) = 3(-1)

OpenStudy (the_fizicx99):

Yesh :3

OpenStudy (the_fizicx99):

Usually its A>0

OpenStudy (anonymous):

Okay, thank you so much!

OpenStudy (the_fizicx99):

no prob yw :}